The sides of a triangular lot are represented by x, 3x, and 3x + 2. Find each side if the perimeter is 632 feet.
The lengths of the sides of the triangle are___ ,____ , and____ .

Answer by colliefan(242)

You already have expressions for the 3 sides. Adding them together gives you the perimeter which is 632 ft. The equation that represents this is:
Then solve for x. Remember, what you do to one side, you must do to the other so it continues to be an equation because each side continues to be equal.

x+3x+3x+2=632
7x+2=632
7x+2-2=632-2
7x+0=630
1/7*7x=630*1/7
1*x=90
x=90
One side is 90 feet. Calculate the other two.
3x = 270 feet
3x+2= 272 feet